OpenStudy (anonymous):

At an amusement park, each premium ride costs $12 and each gold star ride is $8. Stan spends $96 on tickets and he gets 3 more premium rides than gold star rides. How much does he spend on gold star rides?

OpenStudy (anonymous):

@hba can y

OpenStudy (anonymous):

ou help me

OpenStudy (anonymous):

thats not an answer on the test. nice try though :)

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

Is it close or not? :D

OpenStudy (anonymous):

ehhh off by about 20 from the answers on the test

OpenStudy (anonymous):

Is it around $20+? :D

OpenStudy (anonymous):

its 24 lol :)

OpenStudy (anonymous):

Okay, so you don't need the answer anymore. HAHA

OpenStudy (richyw):

\[12P+8G=96\]\[P=G+3\]\[12(G+3)+8G=96\]\[20G=96-36=60\]\[G=3\]so he got three gold ones which costs $24